A lot is in the shape of a right triangle. The shorter leg measures 150m. The hypotenuse is 50m longer than the length of the longer leg. How long is the longer leg?
Answer by nerdybill(7384)

A lot is in the shape of a right triangle. The shorter leg measures 150m. The hypotenuse is 50m longer than the length of the longer leg. How long is the longer leg?
Let x = length (m) of longer leg
then
x+50 = length of hypotenuse
.
applying Pythagorean theorem:
150^2 + x^2 = (x+50)^2
150^2 + x^2 = (x+50)(x+50)
22500 + x^2 = x^2+100x+2500
22500 = 100x+2500
20000 = 100x
200 m = x (longer leg)